There are 64 fifth graders at parks school. 3/8 of them bring their own lunch to school.How many of them bring their own lunch? Show how how you figured out your answer.
Answer by stanbon(75887)

There are 64 fifth graders at parks school. 3/8 of them bring their own lunch to school.How many of them bring their own lunch? Show how how you figured out your answer.
------
Ans: (3/8)64 = 3*8 = 24
===========================
Cheers,
Stan H.